OpenStudy (anonymous):

It costs $2200 to make 100 chairs and $4800 to make 300 chairs. Express the cost as a function of the number of chairs produced, assuming its linear.

OpenStudy (anonymous):

i understand to find the slope you have to do y2-y1/x2-x1

OpenStudy (anonymous):

then its y=13x+ 9

OpenStudy (anonymous):

but then the intercept would be 0, 9

OpenStudy (anonymous):

but you can make 0 chairs and it cost $9

OpenStudy (anonymous):

you cant*

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

m = (4800 - 2200)/(300-100) = 2600/200 = 13 then recall the equation of a line through a point (x1,y1) y-y1 = m(x-x1) y - 4800 = 13(x- 300) y - 4800 = 13x - 3900 y = 13x +900

OpenStudy (anonymous):

but if you make 0 chairs it cant cost $900 right?

OpenStudy (anonymous):

do i just use that equation n put a restriction on it?

OpenStudy (anonymous):

y= 13x+900 when x>0 ?????????

OpenStudy (anonymous):

the maths isn't wrong, i would explain this problem by saying its just a model and models aren't perfect. maybe they make so many chairs usually it doesn't matter what the model says happens at zero

OpenStudy (anonymous):

in the question it says they are making 100 or 300 chairs. maybe there is a minimum chair number we do not know about

OpenStudy (anonymous):

o ok thx!!!!

OpenStudy (anonymous):

happy to help :)